OpenStudy (anonymous):

Help please? Douglas has a segment with endpoints I(5, 2) and J(9, 10) that is divided by a point K such that IK and IJ form a 2:3 ratio. He knows that the distance between the x-coordinates is 4 units. Which of the following fractions will let him find the x-coordinate for point K? 2 over 3 2 over 5 3 over 2 3 over 5

OpenStudy (mrnood):

ok - suppose you had $10 and wanted to give it to 2 people in a ratio of 2:3 how would you divide it?

OpenStudy (anonymous):

Im not sure ive never been good at ratios so i dont know

OpenStudy (mrnood):

Ratios are about fractions To get the fractions you ADD together the 2 numbers in the ratio - that gives you the denominator If you want to divide $20 in a ration of 1:9 then add together the numbers then the fractions are 1/10 and 9/10 to give the ratio so $20 x 1/10 = $2 and $20 x 9/10 = $18 these are in ration 1:9

OpenStudy (mrnood):

so to get the ratio 2:3 write the 2 fractions by adding the numbers for the denominator

OpenStudy (anonymous):

would i add 4 to 2:3?

OpenStudy (mrnood):

Have a look at what I did above the ratio was 1:9 SO I ADDED 1+9 =10 The fractions were then 1/10 & 9/10 so do the same with the ratio 2:3

OpenStudy (anonymous):

So it would be 2/5 and 3/5

OpenStudy (mrnood):

yes - well done one of those is in the answers....

OpenStudy (anonymous):

would it be 2/5?

OpenStudy (mrnood):

are you confident with that? If so - go with it

OpenStudy (anonymous):

I dont knwo. Im never confident with math

OpenStudy (anonymous):

I guess ill go with it then thanks for your help